A pure inductor has a reactance of 22ohmat 200 hertz a.c line .what would be its reactance at 50 hertz line.?

Grade:12

1 Answers

Vikas TU
4 years ago
Hiii 
Reactance of inductor = omega * L
omega = 2*pie*Frequency 
220 = 2*pie*200*L
find L from here and put in another equation Xl (new) = 2*pie*50*L .. this will be your ans.
Hope this helps.